Write an expression for the following (use x for the unknown number). 2 more than 3 times an unknown number is equal to one half of the difference between the unknown number and ten.
Answer by checkley77(12569)

3X+2=(10-X)/2
2(3X+2)=10-X
6X+4=10-X
6X+X=10-4
7X=6
X=6/7
PROOF:
3*6/7+2=(10-6/7)/2
18/7+2=(64/7)*1/2
(18+14)7=64/14
32/7=64/14
32/7=32/7